OpenStudy (anonymous):

help please. algebra word problem. the length of a rectangle is 3 inches more than the width. the area is 18 square inches. find the dimensions.

OpenStudy (shamil98):

You are given. A= 18 A = LW L = 3 + W Plug it in. 18 = (3+W)(W) w^2 + 3w - 18 = 0 (w+6)(w-3) = 0 w=3 (since you cannot have a negative width) A = LW 18 = L(3) 6 = L

OpenStudy (anonymous):

thank you!